### v1

# Towards Robust Multimodal Sentiment Analysis with Incomplete Data

Abstract
The field of Multimodal Sentiment Analysis (MSA) has recently witnessed an emerging direction seeking to tackle the issue of data incompleteness. Recognizing that the language modality typically contains dense sentiment information, we consider it as the dominant modality and present an innovative Language-dominated Noise-resistant Learning Network (LNLN) to achieve robust MSA. The proposed LNLN features a dominant modality correction (DMC) module and dominant modality based multimodal learning (DMML) module, which enhances the model’s robustness across various noise scenarios by ensuring the quality of dominant modality representations. Aside from the methodical design, we perform comprehensive experiments under random data missing scenarios, utilizing diverse and meaningful settings on several popular datasets (e.g., MOSI, MOSEI, and SIMS), providing additional uniformity, transparency, and fairness compared to existing evaluations in the literature. Empirically, LNLN consistently outperforms existing baselines, demonstrating superior performance across these challenging and extensive evaluation metrics. Our code will be available at: https://github.com/Haoyu-ha/LNLN.
1 Introduction
The field of Multimodal Sentiment Analysis (MSA) is at the vanguard of human sentiment identification by assimilating heterogeneous data types, such as video, audio, and language. Its applicability spans numerous fields, prominent among healthcare and human-computer interaction (Jiang et al., 2020). MSA has become essential in enhancing both the precision and the robustness of sentiment analysis by drawing sentiment cues from diverse perspectives.
Changing trends in recent research have taken a turn towards modeling data in natural scenarios from laboratory conditions (Tsai et al., 2019; Hazarika et al., 2020; Yu et al., 2021; Zhang et al., 2023). The shift has created a wider application space in the real-world for MSA, though concerns arise due to problems like sensor failures and problems with Automatic Speech Recognition (ASR), leading to inconsistencies such as incomplete data in real-world deployment.
Numerous impactful solutions have been proposed against this primary concern of incomplete data in multimodal sentiment analysis. For instance, Yuan et al. (2021) introduced a Transformer-based feature reconstruction mechanism, TFR-Net, aiming to strengthen the robustness of the model handling random missing in unaligned multimodal sequences via reconstructing missing data. Furthermore, Yuan et al. (2024) proposed the Noise Intimating-based Adversarial Training (NIAT) model, which superiorly learns a unified joint representation between an original-noisy instance pair, utilizing the attention mechanism and adversarial learning. Lastly, Li et al. (2024) design a Unified Multimodal Missing Modality Self-Distillation Framework (UMDF) which leverages a single network to learn robust inherent representations from consistent multimodal data distributions. Yet, despite these developments, the evaluation metrics for these models are inconsistent, and the evaluation settings are not sufficiently comprehensive. This inconsistency limits effective comparisons and hinders the dissemination of knowledge in the field.
Addressing this gap, our paper aims to offer a comprehensive evaluation on three widely-used datasets, namely MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018) and SIMS (Yu et al., 2020) datasets. We introduce random data missing instances and subsequently compare the performance of existing methods on these datasets. This endeavor seeks to provide an all-encompassing outlook for evaluating the effectiveness and robustness of various methods in the face of incomplete data, thereby sparking new insights in the field.
Additionally, inspired by the previous work ALMT (Zhang et al., 2023), we hypothesize that model robustness improves when the integrity of the dominant modality is preserved despite varying noise levels. Therefore, we introduce a novel model, namely Language-dominated Noise-resistant Learning Network (LNLN), to enhance MSA’s robustness over incomplete data. LNLN aims to augment the integrity of the language modality’s features, regarded as the dominant modality due to its richer sentiment cues, with the support of other auxiliary modalities. The LNLN’s robustness against varying levels of data incompleteness is achieved through a dominant modality correction (DMC) module for dominant modality construction, a dominant modality based multimodal learning (DMML) module for multimodal fusion and classification, and a reconstructor for reconstructing missing information to shield the dominate modality from noise interference. This approach ensures a high-quality dominant modality feature, which significantly bolsters the robustness of LNLN under diverse noise conditions. Consequently, extensive experimental results demonstrate the LNLN’s superior performance across these challenging and extensive evaluation metrics.
In summary, this paper conducts a comprehensive evaluation of existing advanced MSA methods with missing data. This analysis highlights the strengths and weaknesses of various approaches when contending with incomplete data. We believe, it can improve the understanding of different MSA methods’ performance under complex real-world scenarios, thereby informing technology’s future trajectory. Our proposed LNLN also offers valuable insight and guidelines for thriving in this research space.
2 Related Work
2.1 Multimodal Sentiment Analysis
Multimodal Sentiment Analysis (MSA) methods can be categorized into Context-based MSA and Noise-aware MSA, depending on the modeling approach. Most of previous works (Zadeh et al., 2017; Tsai et al., 2019; Mai et al., 2020; Hazarika et al., 2020; Liang et al., 2020; Rahman et al., 2020; Yu et al., 2021; Han et al., 2021; Lv et al., 2021; Yang et al., 2022; Guo et al., 2022; Zhang et al., 2023) can be classified to Context-based MSA. This line of work primarily focuses on learning unified multimodal representations by analyzing contextual relationships within or between modalities. For example, Zadeh et al. (2017) explore computing the relationships between different modalities using the Cartesian product. Tsai et al. (2019) utilize pairs of Transformers to model long dependencies between different modalities. Yu et al. (2021) propose generating pseudo-labels for each modality to further mine the information of consistency and discrepancy between different modalities.
Despite these advances, context-based methods are usually suboptimal under varying levels of noise effects (e.g. random data missing). Several recent works (Mittal et al., 2020; Yuan et al., 2021, 2024; Li et al., 2024) have been proposed to tackle this issue. For example, Mittal et al. (2020) introduce a modality check step to distinguish invalid and valid modalities, achieving higher robustness. Yuan et al. (2024) propose learning a unified joint representation between constructed “original-noisy” instance pairs. Although there have been some advances in improving the model’s robustness under noise scenarios, no extant method has provided a comprehensive and in-depth comparative analysis.
2.2 Robust Representation Learning in MSA
Context-based MSA and Noise-aware MSA differ in their approaches to robust representation learning. In Context-based MSA, robust representation learning typically relies on modeling intra- and inter-modality relationships. For instance, Hazarika et al. (2020) and Yang et al. (2022) apply feature disentanglement to each modality, modeling multimodal representations from multiple feature subspaces and perspectives. Yu et al. (2021) and Liang et al. (2020) explore self-supervised learning and semi-supervised learning to enhance multimodal representations, respectively. Tsai et al. (2019) and Rahman et al. (2020) introduce Transformer to learn the long dependencies of modalities. Zhang et al. (2023) devise a language-guided learning mechanism that uses modalities with more intensive sentiment cues to guide the learning of other modalities.
In contrast, Noise-aware MSA focuses more on perceiving and eliminating the noise present in the data. For example, For example, Mittal et al. (2020) design a modality check module based on metric learning and Canonical Correlation Analysis (CCA) to identify the modality with greater noise. Yuan et al. (2021) design a feature reconstruction network to predict the location of missing information in sequences and reconstruct it. Yuan et al. (2024) introduce adversarial learning (Goodfellow et al., 2014) to perceive and generate cleaner representations.
In our work, the LNLN belongs to the Noise-aware MSA category. Inspired by Zhang et al. (2023), we explore the capability of language-guided mechanisms in resisting noise and aim to provide new perspectives for the study of MSA in noisy scenarios.
3 Method
3.1 Overview
The overall pipeline for the proposed Language-dominated Noise-resistant Learning Network (LNLN) in robust multimodal sentiment analysis is illustrated in Figure 1. As depicted, a crucial initial step involves forming a multimodal input with random data missing, which sets the stage for LNLN training. Once the input is prepared, LNLN first utilizes an embedding layer to standardize the dimension of each modality, ensuring uniformity. Recognizing that language is the dominant modality in MSA (Zhang et al., 2023), a specially designed Dominant Modality Correction (DMC) module employs adversarial learning and a dynamic weighted enhancement strategy to mitigate noise impacts. This module first enhances the quality of the dominant feature computed from the language modality and then integrates them with the auxiliary modalities (visual and audio) in the dominant modality based multimodal learning (DMML) module for effective multimodal fusion and classfication. This process significantly bolsters LNLN’s robustness against various noise levels. Moreover, to refine the network’s capability for fine-grained sentiment analysis, a simple reconstructor is implemented to reconstruct missing data, further enhancing the system’s robustness.
Figure 1: Overall pipeline. Note: $H^{0}_{l}$, $H^{0}_{v}$, $H^{0}_{a}$, $H_{cc}$, and $H^{0}_{p}$ are randomly initialized learnable vectors.
3.2 Input Construction and Multimodal Input
Given the challenges of random data missing, we have constructed data sets that simulate these conditions based on MOSI, MOSEI and SIMS datasets.
Random Data Missing. Following previous protocols (Yuan et al., 2021), for each modality in each sample, we randomly erased changing proportions of information (from 0% to 100%). Specifically, for visual and audio modalities, we fill the erased information with zeros. For language modality, we fill the erased information with [UNK], which indicates the unknown word in BERT (Kenton and Toutanova, 2019).
Multimdoal Input. For each sample in the dataset, we incorporate data from three modalities: language, audio, and visual data. Consistent with previous works (Zhang et al., 2023), each modality is processed using widely-used tools: language data is encoded using BERT (Kenton and Toutanova, 2019), audio features are extracted through Librosa (McFee et al., 2015), and visual features are obtained using OpenFace (Baltrusaitis et al., 2018). These pre-processed inputs are represented as sequences, denoted by $U^{0}_{m}\in\mathbb{R}^{T_{m}\times d_{m}}$, where $m\in\{l,v,a\}$ represents the modality type ($l$ for language, $v$ for visual, $a$ for audio), $T_{m}$ indicates the sequence length, and $d_{m}$ refers to the dimension of each modality’s vector. With obtained $U^{0}_{m}$, we apply random data missing to $U^{0}_{m}$, thus forming the noise corrupted multimodal input $U^{1}_{m}\in\mathbb{R}^{T_{m}\times d_{m}}$.
3.3 Dominant Modality based Multimodal Learning
Inspired by previous work ALMT, we hypothesize that model robustness improves when the integrity of the dominant modality is preserved despite varying noise levels. We improve ALMT based on the designed DMC module and Reconstructor, thus implementing dominant modality based DMML module for sentiment analysis under random data missing scenarios. Here, we mainly introduce the parts that differ from ALMT. Further details available in Zhang et al. (2023).
Modality Embedding. For multimodal input $U^{1}_{m}$, we employ an Embedding Encoder $\operatorname{E}^{1}_{m}$ with two Transformer encoder layers to extract and unify the feature. Each modality begins with a randomly initialized low-dimensional token $H_{m}^{0}\in\mathbb{R}^{T\times d_{m}}$. These tokens are then processed by the Transformer encoder layer, embedding essential modality information and producing unified features, represented as $H_{m}^{1}\in\mathbb{R}^{T\times d}$. The process is formalized by:
$H_{m}^{1}=\operatorname{E}^{1}_{m}\left(\operatorname{concat}\left(H_{m}^{0},U^{1}_{m}\right)\right),$ (1)
where $\operatorname{E}^{1}_{m}(\cdot)$ extracts features for each modality, and $\operatorname{concat}(\cdot)$ represents the concatenation operation.
Adaptive Hyper-modality Learning. In the original ALMT, each Adaptive Hyper-modality Learning layer contains a Transformer and two multi-head attention (MHA) modules. These are applied to learn language representations at different scales and hyper-modality representations from visual and audio modalities, guided by the language modality. Considering the possibility of severe interference in the language modality (i.e. dominant modality) due to random data missing, we designed a Dominant Modality Correction (DMC) module to generate the proxy dominant feature $H^{1}_{p}$ and construct corrected dominate feature $H^{1}_{d}$ (more details can be found in Section 3.4). Specifically, the process of learning corrected dominated representation $H_{d}^{i}$ at different scales can be described as:
$H_{d}^{i}=\operatorname{E}^{i}_{m}\left(H_{d}^{i-1}\right),$ (2)
where $i\in\{2,3\}$ means the $i$-th layer of Adaptive Hyper-modality Learning module, $\operatorname{E}^{i}_{m}(\cdot)$ is the i-th Transformer encoder layer, $H_{d}^{i}\in\mathbb{R}^{T\times d}$ is corrected dominated feature at different scale. To learn the hyper-modality representation, the corrected dominated feature and audio/visual features are used to calculate Query and Key/Value, respectively. Briefly, the process can be written as follows:
$H_{\text{hyper}}^{i}=H_{\text{hyper }}^{i-1}+\operatorname{MHA}(H_{d}^{i},H_{a}^{1})+\operatorname{MHA}(H_{d}^{i},H_{v}^{1}),$ (3)
where $\operatorname{MHA}(\cdot)$ represents multi-head attention, $H_{\text{hyper}}^{i}\in\mathbb{R}^{T\times d}$ is the hyper-modality feature. Note that the feature $H_{\text{hyper}}^{0}\in\mathbb{R}^{T\times d}$ is a random initialized vector.
Multimodal Fusion and Prediction. With obtained $H_{d}^{3}$ and $H_{hyper}^{3}$, a Transformer encoder with a classifier at a depth of 4 layers is employed for multimodal fusion and sentiment prediction:
$\hat{y}=\operatorname{CrossTransformer}(H_{d}^{3},H^{3}_{hyper}),$ (4)
where $\hat{y}$ is the sentiment prediction.
3.4 Dominant Modality Correction
This module consists of two steps, i.e., completeness check of dominant modality and proxy dominant feature generation using adversarial learning (Ganin and Lempitsky, 2015; Goodfellow et al., 2014).
Completeness Check. We apply an encoder $\operatorname{E}_{cc}$ that consists of a Transformer encoder with the depth of two layers and a classifier for completeness check. For example, if the missing rate of dominate modality is 0.3, the label of completeness is 0.7. This completeness prediction $w$ can be obtained as follows:
$w={\operatorname{E}_{cc}\left(\operatorname{Concat}\left(H_{cc},H_{l}^{1}\right)\right)},$ (5)
where $H_{cc}\in\mathbb{R}^{T\times d}$ is a randomly initialized token for completeness prediction. We optimize this process using L2 loss:
$\mathcal{L}_{cc}=\frac{1}{N_{b}}\sum_{k=0}^{N_{b}}\left\|w^{k}-\hat{w}^{k}\right\|_{2}^{2},$ (6)
where $N_{b}$ is the number of samples in the training set, $\hat{w}^{k}$ is the label of completeness of $k$-th sample.
Proxy Dominant Feature Generation. With the randomly initialized feature $H^{0}_{p}\in\mathbb{R}^{T\times d}$, visual feature $H^{1}_{v}$ and audio feature $H^{1}_{a}$, we employ a Proxy Dominant Feature Generator $E_{DFG}$, which consists of two Transformer encoder layers. This setup generates the proxy dominant features $H^{1}_{p}\in\mathbb{R}^{T\times d}$, designed to complement and correct the dominant modality. The corrected dominant feature $H^{1}_{d}\in\mathbb{R}^{T\times d}$ is calculated by combining $H^{1}_{p}$ and the language feature $H^{1}_{l}$, weighted by the predicted completeness $w$:
$H^{1}_{p}=\operatorname{E_{DFG}}\left(\operatorname{Concat}\left(H^{0}_{p},H_{a}^{1},H_{v}^{1}\right),\theta_{DFG}\right),$ (7)
$H^{1}_{d}=(1-w)*H^{1}_{p}+w*H^{1}_{l},$ (8)
where $\theta_{DFG}$ denotes the parameters of the Proxy Dominant Feature Generator $E_{DFG}$.
To ensure that the agent feature offers a distinct perspective from the visual and audio features, we utilize an effectiveness discriminator $D$. This discriminator includes a binary classifier and a Gradient Reverse Layer (GRL) (Ganin and Lempitsky, 2015) and is tasked with identifying the origin of the agent features:
$\hat{y}_{p}=\operatorname{D}\left(H^{1}_{p}/H^{1}_{l},\theta_{D}\right),$ (9)
where $\theta_{D}$ represents the parameters of the effectiveness discriminator $D$, and $\hat{y}_{p}$ indicates the prediction of whether the input feature originates from the language modality.
In practice, the generator and the discriminator engage in an adversarial learning structure. The discriminator aims to identify whether the features are derived from the language modality, while the generator’s objective is to challenge the discriminator’s ability to make accurate predictions. This dynamic is encapsulated in the adversarial learning objective:
$\min_{\theta_{D}}\max_{\theta_{DFG}}\mathcal{L}_{adv}=-\frac{1}{N_{b}}\sum_{k=0}^{N_{b}}y^{k}_{p}\cdot\log\hat{y}^{k}_{p},$ (10)
where $N_{b}$ is the number of samples in the training set, and $y^{k}_{p}$ indicates the label determining whether the input feature for the $k$-th sample originates from the visual or audio modality.
3.5 Reconstructor
Our experiments demonstrate that reconstructing missing information can significantly enhance regression metrics. More details about this are shown in Table 3. To address this, we have developed a reconstructor, denoted as $E_{rec}$, which comprises two Transformer layers designed to effectively rebuild missing information of each modality. The operational equation for the reconstructor is:
$\hat{U}^{0}_{m}={\operatorname{E}_{m}^{rec}\left(U^{1}_{m}\right)}$ (11)
where $\hat{U}^{0}_{m}$ is the reconstructed feature corresponding to the feature $U^{0}_{m}$.
To optimize the performance of the reconstructor, we apply an L2 loss function:
$\mathcal{L}_{rec}=\frac{1}{N_{b}}\sum_{h=0}^{N_{b}}\sum_{m}\left\|U_{m}^{0}{}^{k}-\hat{U}_{m}^{0}{}^{k}\right\|_{2}^{2},$ (12)
where $U_{m}^{0}{}^{k}$ and $\hat{U}_{m}^{0}{}^{k}$ represent the original and reconstructed features with missing information for the $k$-th sample, respectively. This loss function helps minimize the discrepancies between the original and reconstructed features, thereby improving the accuracy of other components, such as Dominant Modality Correction and final sentiment prediction.
3.6 Overall Learning Objectives
To sum up, our method involves four learning objectives, including a completeness check loss $\mathcal{L}_{cc}$, an adversarial learning loss $\mathcal{L}_{adv}$ for proxy dominant feature generation, a reconstruction loss $\mathcal{L}_{rec}$ and one final sentiment prediction loss $\mathcal{L}_{sp}$. The sentiment prediction loss $\mathcal{L}_{sp}$ can be described as:
$\mathcal{L}_{sp}=\frac{1}{N_{b}}\sum_{n=0}^{N_{b}}\left\|y^{n}-\hat{y}^{n}\right\|_{2}^{2},$ (13)
Therefore, the overall loss $\mathcal{L}$ can be written as:
$\mathcal{L}=\alpha\mathcal{L}_{cc}+\beta\mathcal{L}_{adv}+\gamma\mathcal{L}_{rec}+\delta\mathcal{L}_{sp},$ (14)
where $\alpha$, $\beta$, $\gamma$ and $\delta$ are hyper-parameters. On MOSI and MOSEI datasets, we empirically set them to 0.9, 0.8, 0.1, and 1.0, respectively. On SIMS dataset, we empirically set them to 0.9, 0.6, 0.1, and 1.0, respectively.
4 Experiments and Analysis
In this section, we provide a comprehensive and fair comparison between the proposed LNLN and previous state-of-the-art MSA methods on MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018) and SIMS (Yu et al., 2020) datasets. 1) For each method, we select three random seeds for training and evaluation. 2) We report the performance of the model on a variety of evaluation metrics, which is much more comprehensive compared to some of the previous work.
4.1 Datasets
MOSI. The dataset includes 2,199 multimodal samples, integrating visual, audio, and language modalities. It is divided into a training set of 1,284 samples, a validation set of 229 samples, and a test set of 686 samples. Every single sample has been given a sentiment score, varying from -3, indicating strongly negative sentiment, to 3, signifying strongly positive sentiment.
MOSEI. The dataset consists of 22,856 video clips sourced from YouTube. The sample is divided into 16,326 clips for training, 1,871 for validation, and 4,659 for testing. Each clip is carefully labeled with a sentiment score, ranging from -3, representing the strongly negative sentiment, to 3, denoting the strongly positive sentiment.
SIMS. The dataset is a Chinese multimodal sentiment dataset that includes 2,281 video clips sourced from different movies and TV series. It has been partitioned into 1,368 samples for training, 456 for validation, and 457 for testing. Each sample has been manually annotated with a sentiment score ranging from -1 (negative) to 1 (positive).
4.2 Evaluation Settings and Criteria
For a fair and comprehensive evaluation, we experiment ten times, setting the missing rates $r$ to predefined values from 0 to 0.9 with an increment of 0.1. For instance, 50% of the information is randomly erased from each modality in the test data when $r=0.5$. Unlike previous works (Yuan et al., 2021, 2024), we did not evaluate at $r=1.0$, as this would imply complete data erasure from each modality, rendering the experiment non-informative. With the obtained results of each missing rate, we compute the average value as the model’s overall performance under different levels of noise.
For evaluation criteria, we report the binary classification accuracy (Acc-2), the F1 score associated with Acc-2, and the mean absolute error (MAE). For Acc-2, we calculated accuracy and F1 in two ways: negative/non-negative and negative/positive on the MOSI and MOSEI datasets, respectively. Additionally, we provide the three-class accuracy (Acc-3), the seven-class accuracy (Acc-7), and the correlation of the model’s prediction with humans (Corr) on the MOSI dataset. For the SIMS dataset, we report Acc-3, the five-class accuracy (Acc-5), and Corr.
4.3 Robustness Comparison
Tables 1 and 2 show the robustness evaluation results on the MOSI, MOSEI, and SIMS datasets. As shown in Table 1, LNLN achieves state-of-the-art performance on most metrics. For example, on the MOSI dataset, LNLN achieved a relative improvement of 9.46% on Acc-7 compared to the sub-optimal result obtained by MMIM, demonstrating the robustness of LNLN in the face of different noise effects. However, on the MOSEI dataset, LNLN achieves only sub-optimal performance on metrics such as Acc-7 and Acc-5. After analyzing the data distribution (see Appendix B.3) and the confusion matrix (see Appendix B.5), we believe that this is due to the fact that LNLN does not overly bias the neutral category, which has a disproportionate number of samples in noisy scenarios.
As shown in Table 2, LNLN achieves a remarkable improvement in F1 on the SIMS dataset, with a relative improvement of 4.23% on F1 compared to the sub-optimal result obtained by ALMT. Similar to CENET’s performance on the MOSEI dataset, TETFN on the SIMS dataset has a strong tendency to predict inputs as weakly negative categories with high sample sizes, resulting in seemingly better performance on some metrics.
Additionally, as shown in Figure 2, we present the performance curves of several advanced methods under varying missing rates. The results demonstrate that the proposed LNLN consistently outperforms others across most scenarios, showing its robustness under different missing rates.
Overall, LNLN attempts to make predictions in the face of highly noisy inputs without the severe lazy behavior observed in other models, which often leads to predictions heavily biased towards a certain category. This demonstrates that LNLN shows strong robustness and competitiveness across various datasets and noise levels, highlighting its effectiveness in multimodal sentiment analysis.
Table 1: Robustness comparison of the overall performance on MOSI and MOSEI datasets. Note: The smaller MAE indicate the better performance.
Method MOSI MOSEI
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-7 Acc-5 Acc-2 F1 MAE Corr
MISA 29.85 33.08 71.49 / 70.33 71.28 / 70.00 1.085 0.524 40.84 39.39 71.27 / 75.82 63.85 / 68.73 0.780 0.503
Self-MM 29.55 34.67 70.51 / 69.26 66.60 / 67.54 1.070 0.512 44.70 45.38 73.89 / 77.42 68.92 / 72.31 0.695 0.498
MMIM 31.30 33.77 69.14 / 67.06 66.65 / 64.04 1.077 0.507 40.75 41.74 73.32 / 75.89 68.72 / 70.32 0.739 0.489
CENET 30.38 37.25 71.46 / 67.73 68.41 / 64.85 1.080 0.504 47.18 47.83 74.67 / 77.34 70.68 / 74.08 0.685 0.535
TETFN 30.30 34.34 69.76 / 67.68 65.69 / 63.29 1.087 0.507 30.30 47.70 69.76 / 67.68 65.69 / 63.29 1.087 0.508
TFR-Net 29.54 34.67 68.15 / 66.35 61.73 / 60.06 1.200 0.459 46.83 34.67 73.62 / 77.23 68.80 / 71.99 0.697 0.489
ALMT 30.30 33.42 70.40 / 68.39 72.57 / 71.80 1.083 0.498 40.92 41.64 76.64 / 77.54 77.14 / 78.03 0.674 0.481
LNLN 34.26 38.27 72.55 / 70.94 72.73 / 71.25 1.046 0.527 45.42 46.17 76.30 / 78.19 77.77 / 79.95 0.692 0.530
Table 2: Robustness comparison of the overall performance on SIMS dataset. Note: The smaller MAE indicate the better performance.
Method Acc-5 Acc-3 Acc-2 F1 MAE Corr
MISA 31.53 56.87 72.71 66.30 0.539 0.348
Self-MM 32.28 56.75 72.81 68.43 0.508 0.376
MMIM 31.81 52.76 69.86 66.21 0.544 0.339
CENET 22.29 53.17 68.13 57.90 0.589 0.107
TETFN 33.42 56.91 73.58 68.67 0.505 0.387
TFR-Net 26.52 52.89 68.13 58.70 0.661 0.169
ALMT 34.16 56.47 71.85 76.21 0.509 0.372
LNLN 34.64 57.14 72.73 79.43 0.514 0.397
Figure 2: Performance curves of various missing rates. (a), (b) and (c) is the F1 curves on MOSI, MOSEI and SIMS, respectively. (d), (e) and (f) is the MAE curves on MOSI, MOSEI and SIMS, respectively. Note: The smaller MAE indicate the better performance.
4.4 Effects of Different Components
To verify the role of different modules, we present the ablation results after subtracting each component on the MOSI dataset separately. As shown in Table 3, a few metrics show an upward trend in performance when certain modules are removed. On one hand, we believe this is due to randomness in noisy scenarios. On the other hand, the same hyper-parameters cannot make the model perform optimally across all metrics.
Moreover, we can see that the performance of LNLN decreases relatively slightly when the DMC and reconstructor are removed individually. However, there is a significant drop in performance when both are removed, e.g., 4.72% decrease in Acc-7 and 5.87% decrease in Acc-5. Performance drops further after replacing DMML with concatenation fusion, which also proves that DMML plays an important role in MSA. Additionally, after removing the noisy data used for training, the performance of LNLN also shows a significant decrease. This observation indicates that LNLN needs to learn to be aware of noise with the help of noisy data.
Overall, both the presence of noisy data and well-designed components are crucial for robust MSA in scenarios with random missing data.
Table 3: Effects of different components. Note: The smaller MAE indicate the better performance.
Method MOSI SIMS
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-5 Acc-3 Acc-2 F1 MAE Corr
LNLN 34.26 38.27 72.55 / 70.94 72.73 / 71.25 1.046 0.527 34.64 57.14 72.73 79.43 0.514 0.397
w/o DMC & Reconstructor 29.54 32.40 71.98 / 70.31 73.53 / 71.29 1.104 0.478 33.81 54.00 72.57 78.10 0.525 0.376
w/o DMC 33.94 37.89 72.24 / 71.02 72.86 / 71.45 1.049 0.522 33.96 58.30 73.40 77.58 0.518 0.386
w/o Reconstructor 33.94 38.52 70.81 / 71.35 71.17 / 71.69 1.054 0.518 34.02 58.29 73.07 78.71 0.548 0.351
w/o DMML 33.22 37.25 69.55 / 69.55 70.81 / 70.71 1.084 0.504 34.87 56.06 70.74 76.94 0.553 0.245
w/o Missing Data for Training 30.11 34.82 70.62 / 68.78 72.49 / 71.03 1.105 0.486 33.48 53.09 72.96 79.17 0.519 0.360
4.5 Effects of Different Regularization
As shown in Table 4, we present the ablation results after subtracting each regularization on the MOSI dataset separately. Obviously, after each regularization is removed, LNLN shows a decrease in performance on most metrics. This demonstrates that regularization contributes to the learning of the model. Similar to the results in Table 3, there are a few metrics that improve after the removal of regularization. We believe this is mainly due to the difficulty of optimizing a single set of hyper-parameters to achieve the best performance across all metrics.
Table 4: Effects of different regularization. Note: The smaller MAE indicate the better performance.
Method MOSI SIMS
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-5 Acc-3 Acc-2 F1 MAE Corr
LNLN 34.26 38.27 72.55 / 70.94 72.73 / 71.25 1.046 0.527 34.64 57.14 72.73 79.43 0.514 0.397
w/o $\mathcal{L}_{cc}$ 33.61 37.78 72.32 / 70.72 72.55 / 71.02 1.047 0.531 32.87 56.41 70.78 75.30 0.523 0.376
w/o $\mathcal{L}_{adv}$ 32.84 37.79 72.28 / 70.31 72.83 / 70.40 1.052 0.526 32.47 54.24 73.80 73.91 0.525 0.351
w/o $\mathcal{L}_{rec}$ 33.94 38.52 72.46 / 70.81 72.69 / 71.17 1.048 0.523 34.33 56.12 71.29 74.73 0.525 0.364
4.6 Effects of Different Modalities
To verify the role of different modalities, we present the ablation results after subtracting each modality from the MOSI dataset. As shown in Table 5, all modalities contribute to the final performance. Notably, when the language modality is removed, the performance of LNLN shows a significant decrease, while removing the visual and audio modalities results in a smaller decrease. This indicates that the language modality plays a crucial role in the MSA dataset.
Moreover, we observe that LNLN converges on most metrics even when the visual and audio modalities are removed. This is due to the model’s lazy behavior, where LNLN tends to fix its predictions to a certain categories. More details can be found in Appendix B.3 and B.5.
Table 5: Effects of different modalities. Note: The smaller MAE indicate the better performance.
Method MOSI
Acc-7 Acc-5 Acc-2 F1 MAE Corr
LNLN 34.26 38.27 72.55 / 70.94 72.73 / 71.25 1.046 0.527
w/o language 16.30 16.31 54.88 / 55.79 60.33 / 63.10 1.399 0.033
w/o audio 33.51 38.10 72.31 / 70.54 73.67 / 71.10 1.064 0.535
w/o vision 33.53 38.15 72.32 / 70.58 73.68 / 71.14 1.064 0.536
4.7 Case Study
As shown in Figure 3, we visualize several successful and failed predictions made by LNLN and ALMT from MOSI dataset for case study. It shows that LNLN can perceives sentiment cues in challenging samples, which demonstrates its ability to capture sentiment cues in noisy scenario. However, for inputs with high missing rates (e.g., the third example in the figure), both LNLN and ALMT fail to make correct predictions. We believe this is due to the high loss of valid information in the multimodal input, making accurate predictions difficult.
Figure 3: Visualization of successful and failure predictions.
4.8 Discussion
In this work, we conduct a comprehensive evaluation of current advanced methods, the results of which have triggered some thoughts. Specifically, we find that all methods fail to accurately predict the inputs in high missing rate scenarios because most of them are designed specifically for complete data (see Appendix B.5 for more details). However, some methods show relatively more robust performance in low missing rate scenarios and some specific scenarios (see Appendix B.7), such as Self-MM, TETFN, and ALMT. We believe that some of the ideas in these methods are useful for future research on more robust MSA for incomplete data.
For example, the Unimodal Label Generation Module (ULGM) (Yu et al., 2021) used in both Self-MM and TETFN facilitates the modeling of incomplete data. Due to the random data missing, the temporal and structured affective information in the data is corrupted, making it difficult for the models to perceive consistency and variability information between different modalities. Using ULGM to generate pseudo-labels as an additional supervisory signal for the models may directly enable the model to learn the relationship between different modalities. For ALMT, we believe that its Adaptive Hyper-modality Learning (AHL) module facilitates the modeling of robust affective representations. AHL mainly uses the dominant modality to compute query vectors and applies multi-head attention to query useful sentiment cues from other modalities as a complement to the dominant modality. This process may reduce the difficulty of multimodal fusion by avoiding the introduction of sentiment-irrelevant information to some extent. Our proposed LNLN, based on this idea from ALMT, proves the effectiveness of the method.
Additionally, we explore the performance of these methods in modality missing scenarios, which is a special case of random data missing where the partial modality missing rate $r=1.0$ (see Appendix B.7 for more details). We find that ALMT and LNLN demonstrate advantages over other methods in terms of F1 metrics when the language modality is completely lost.
5 Conclusion and Future Work
In this paper, a novel Language-dominated Noise-resistant Learning Network (LNLN) is proposed to better learn sentiment cues for robust MSA. Due to the collaboration between Dominate Modality Correction (DMC) module and Dominant Modality based Multimodal Learning (DMML) module, and Reconstructor, LNLN can dynamically enhancing the dominant modality to achieve superior performance in data missing scenarios with varying levels. Extensive evaluation shows that none of the existing methods can effectively model the data with high missing rates. The related analyses can provide suggestions for other researchers to better handle robust MSA. In the future, we will focus on improving the generalization of the model to handle different types of scenes and varying intensities of noise effectively.
Limitations
We believe there are several limitations to the LNLN: 1) The LNLN achieves well performance in data missing scenarios, but not always better than all other methods in the modality missing scenarios, which demonstrates the lack of multi-scene generalization capability of the LNLN. 2) The data in real-world scenarios is much more complex. In addition to the presence of missing data, other factors need to be considered, such as diverse cultural contexts, varying user behavior patterns, and the influence of platform-specific features on the data. These factors can introduce additional noise and variability, which may require further model adaptation and tuning to handle effectively. 3) Tuning the hyperparameters, particularly those related to the loss functions, can be challenging and may require more sophisticated methods to achieve optimal performance.
Appendix A Implementation Details
We used PyTorch 2.2.1 to implement the method. The experiments were conducted on a PC with an AMD EPYC 7513 CPU at 2.6GHz and 1.0TB memory, and an NVIDIA Tesla A40 with 48GB memory. To ensure consistent and fair comparisons across all methods, we conducted each experiment three times using fixed random seeds of $1$, 1112, and 1113, respectively. Details of the key parameters used in the experiments are shown in Table 6.
In addition, the result of MISA, Self-MM, MMIM, CENET, TETFN, and TFR-Net is reproduced by the authors from open source code in the MMSA [^1] MMSA: https://github.com/thuiar/MMSA (Mao et al., 2022), which is a unified framework for MSA, using default hyper-parameters. The result of ALMT is reproduced by the authors from open source code on Github [^2] ALMT: https://github.com/Haoyu-ha/ALMT using default hyper-parameters.
Table 6: Hyper-parameters of LNLN we use on the different datasets
MOSI MOSEI SIMS
Vector Length $T$ 8 8 8
Vector Dimension $d$ 128 128 128
Batch Size 64 64 64
Initial Learning Rate 1e-4 1e-4 1e-4
Loss Weight $\alpha$, $\beta$, $\gamma$, $\delta$ 0.9, 0.8, 0.1, 1 .0 0.9, 0.8, 0.1, 1.0 0.9, 0.6, 0.1, 1.0
Optimizer AdamW AdamW AdamW
Epochs 200 200 200
Warm Up
Cosine Annealing
Early Stop
Seed $1$,1112,1113 $1$,1112,1113 $1$,1112,1113
Appendix B Additional Experiments and Analysis
B.1 Effect of Regularization Weight on Model Performance
As shown in Table 7, we empirically tried different combinations of hyperparameters in the loss function. The results shows that the selected parameters can make the LNLN achieve the better performance in most metrics. This also demonstrates the effectiveness of the optimization objectives of LNLN.
Table 7: Effect of Regularization Weight on Model Performance. Note: The smaller MAE indicate the better performance.
MOSI
$\alpha$, $\beta$, $\gamma$, $\delta$ Acc-7 Acc-5 Acc-2 F1 MAE Corr
0.9, 0.8, 0.1, 1.0 34.26 38.27 72.55 / 70.94 72.73 / 71.25 1.046 0.527
0.9, 0.8, 0, 1.0 33.94 38.52 72.46 / 70.81 72.69 / 71.17 1.148 0.523
0.9, 0.8, 0.2, 1.0 32.76 37.33 72.20 / 70.69 72.43 / 71.22 1.078 0.511
0.9, 0, 0.1, 1.0 32.84 37.79 72.28 / 70.31 72.83 / 70.40 1.052 0.526
0.9, 0.5, 0.1, 1.0 32.43 37.23 72.04 / 70.24 72.59 / 70.96 1.063 0.511
0.9, 1.0, 0.1, 1.0 32.24 37.07 71.61 / 70.72 73.04 / 70.89 1.063 0.518
0, 0.8, 0.1, 1.0 33.61 37.78 72.32 / 70.72 72.55 / 71.02 1.047 0.531
0.5, 0.8, 0.1, 1.0 33.29 37.78 72.40 / 70.95 72.61 / 71.29 1.067 0.515
1.0, 0.8, 0.1, 1.0 33.22 36.64 71.66 / 70.57 72.94 / 71.80 1.067 0.511
SIMS
$\alpha$, $\beta$, $\gamma$, $\delta$ Acc-5 Acc-3 Acc-2 F1 MAE Corr
0.9, 0.6, 0.1, 1.0 34.64 57.14 72.73 79.43 0.514 0.397
0.9, 0.6, 0, 1.0 32.87 56.41 70.78 75.30 0.523 0.376
0.9, 0.6, 0.2, 1.0 33.95 55.11 70.82 73.20 0.517 0.390
0.9, 0, 0.1, 1.0 32.47 54.24 73.80 73.91 0.525 0.351
0.9, 0.3, 0.1, 1.0 35.16 56.57 72.86 77.54 0.524 0.358
0.9, 1.0, 0.1, 1.0 32.92 54.66 72.50 76.94 0.543 0.363
0, 0.6, 0.1, 1.0 34.33 56.12 71.29 74.73 0.525 0.364
0.5, 0.6, 0.1, 1.0 33.96 55.57 70.63 73.06 0.539 0.360
1.0, 0.6, 0.1, 1.0 33.25 57.03 70.57 72.20 0.522 0.367
B.2 Analysis of Model Stability
To evaluate the stability of the model, we selected the MOSI dataset for ablation experiments. Specifically, based on the overall performance in Table 1, we selected several representative methods to additionally compute the standard deviation. For each method, we first calculate the standard deviation of evaluation results for the three random seeds when the missing rate $r$ ranges from 0 to 0.9, and then average the 10 sets of standard deviation. The final result is shown in Table 8. We can see that the standard deviation of these methods is not very large in most metrics, indicating that all these methods can guarantee stability in the presence of random noise. It should be noted that LNLN strikes a balance between overall performance and stability.
Table 8: Comparison of Model Stability on MOSI dataset. Note: -$\pm$- means mean±$\pm$±std. The smaller MAE indicate the better performance.
Method Acc-7 Acc-5 Acc-2 F1 MAE Corr
MISA 29.85$\pm$2.62 33.08$\pm$1.77 71.49$\pm$1.00 / 70.33$\pm$0.93 71.28$\pm$0.92/ 70.00$\pm$1.26 1.085$\pm$0.08 0.524$\pm$0.08
Self-MM 29.55$\pm$1.06 34.67$\pm$1.87 70.51$\pm$0.71 / 69.26$\pm$1.07 66.60$\pm$1.92 / 67.54$\pm$2.34 1.070$\pm$0.13 0.512$\pm$0.07
CENET 30.38$\pm$1.27 37.25$\pm$1.53 71.46$\pm$0.60 / 67.73$\pm$0.71 68.41$\pm$1.12 / 64.85$\pm$2.56 1.080$\pm$0.14 0.504$\pm$0.11
TFR-Net 29.54$\pm$1.00 34.67$\pm$1.75 68.15$\pm$1.25 / 66.35$\pm$1.09 61.73$\pm$2.82 / 60.06$\pm$2.33 1.200$\pm$0.11 0.459$\pm$0.25
LNLN 34.26$\pm$1.17 38.27$\pm$1.23 72.55$\pm$1.11 / 70.94$\pm$1.28 72.73$\pm$0.99 / 71.25$\pm$1.06 1.046$\pm$0.21 0.527$\pm$0.17
B.3 Analysis of Data Distribution
Figure 4 illustrates the data distribution on the MOSI, MOSEI, and SIMS datasets. Significant category imbalances can be seen across all datasets. Additionally, the distributions of the training set, validation set, and test set are not identical on the MOSI and MOSEI datasets. For example, as shown in Figure 4 (a), the percentage of weakly negative samples is 15.5% in the MOSI training set, while this category reaches 22.7% in the test set, representing the highest percentage among all categories. Similarly, Figure 4 (b) and Figure 4 (c) show that this phenomenon also exists in the MOSEI and SIMS datasets.
The unbalanced data distribution makes it difficult for the model to perform well in data missing scenarios. Specifically, it is easy for the model to engage in lazy behavior in high-noise scenarios, simply biasing predictions towards categories with a higher proportion in the training set. These unbalanced distributions may further increase the learning difficulty of the model when facing data missing. More details can be found in Appendix B.5.
B.4 Details of Robust Comparison
Tables 9 and 10 show the details of robust comparison on the MOSI, MOSEI, and SIMS datasets, respectively. We observe that Self-MM and TETFN have a clear advantage in many evaluation metrics when the missing rate $r$ is low. As the missing rate $r$ increases, LNLN shows a significant improvement in all evaluation metrics. This demonstrates LNLN’s ability to effectively model data affected by noise of varying intensity. In general, models augmented with noisy data in training usually cannot achieve state-of-the-art performance when the noise is low due to differences in data distribution. It is also worthwhile to study in the future how to balance the performance of models under different levels of noise.
In addition, when $r$ is high, many models converge in their performance on metrics such as Acc-7 and Acc-5 (see Appendix B.5). This is due to the fact that these models exhibit lazy behavior, tending to predict inputs as categories with a higher number of samples in the training set. In such cases, we believe that the models do not actually learn effective knowledge, even if their performance appears good.
Figure 4: Data distribution of MOSI, MOSEI and SIMS datasets.
Table 9: Details of Robust Comparison on MOSI and MOSEI with different random missing rates. Note: The smaller MAE indicate the better performance. Note: The smaller MAE indicate the better performance.
Method MOSI MOSEI
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-7 Acc-5 Acc-2 F1 MAE Corr
Random Missing Rate $r=0$
MISA 43.05 48.30 82.78 / 81.24 82.83 / 81.23 0.771 0.777 51.79 53.85 85.28 / 84.10 85.10 / 83.75 0.552 0.759
Self-MM 42.81 52.38 85.22 / 83.24 85.19 / 83.26 0.720 0.790 53.89 55.72 85.34 / 84.68 85.11 / 84.66 0.531 0.764
MMIM 45.92 49.85 83.43 / 81.97 83.43 / 81.94 0.744 0.778 50.76 53.04 83.53 / 81.65 83.39 / 81.41 0.576 0.724
CENET 43.20 50.39 83.08 / 81.49 83.06 / 81.48 0.748 0.785 54.39 56.12 85.49 / 82.30 85.41 / 82.60 0.531 0.770
TETFN 44.07 51.31 82.62 / 81.10 82.67 / 81.09 0.719 0.794 44.07 55.96 82.62 / 81.10 82.67 / 81.09 0.719 0.794
TFR-Net 40.82 47.91 83.64 / 81.68 83.57 / 81.61 0.805 0.760 53.71 47.91 84.96 / 84.65 84.71 / 84.34 0.550 0.745
ALMT 42.37 48.49 84.91 / 82.75 85.01 / 82.94 0.752 0.768 52.18 53.89 85.62 / 83.99 85.69 / 84.53 0.542 0.752
LNLN 44.56 49.76 84.25 / 81.24 84.61 / 81.79 0.751 0.778 50.66 51.94 84.14 / 83.61 84.53 / 84.02 0.572 0.735
Random Missing Rate $r=0.1$
MISA 40.28 46.21 80.18 / 79.01 80.21 / 78.97 0.847 0.721 50.13 51.34 82.21 / 82.28 81.28 / 80.79 0.598 0.722
Self-MM 40.33 49.03 81.40 / 80.03 81.19 / 80.03 0.812 0.728 51.80 53.18 83.03 / 83.79 82.43 / 83.23 0.564 0.725
MMIM 42.61 46.65 79.98 / 78.13 79.83 / 77.99 0.825 0.718 49.09 51.19 82.00 / 81.09 81.57 / 80.15 0.602 0.696
CENET 40.13 46.60 80.08 / 78.38 79.91 / 78.20 0.837 0.719 52.83 54.23 83.75 / 82.41 83.42 / 82.34 0.556 0.739
TETFN 40.67 46.84 80.59 / 78.91 80.55 / 78.79 0.805 0.731 40.67 54.28 80.59 / 78.91 80.55 / 78.79 0.805 0.731
TFR-Net 38.63 45.82 79.27 / 77.99 78.70 / 77.61 0.872 0.705 52.29 45.82 82.92 / 83.31 82.25 / 82.40 0.573 0.715
ALMT 39.84 45.48 80.90 / 78.67 81.15 / 79.08 0.843 0.703 49.98 51.38 84.14 / 82.84 84.23 / 83.04 0.583 0.718
LNLN 42.37 47.91 81.20 / 78.43 81.62 / 79.04 0.820 0.724 49.96 51.25 83.32 / 82.73 83.66 / 82.91 0.591 0.712
Random Missing Rate $r=0.2$
MISA 36.25 41.55 77.54 / 76.34 77.58 / 76.30 0.939 0.654 47.24 47.66 77.84 / 79.93 75.56 / 76.88 0.659 0.674
Self-MM 36.64 43.98 78.15 / 76.48 77.76 / 76.51 0.901 0.660 49.44 50.51 80.84 / 82.33 79.76 / 81.17 0.604 0.678
MMIM 39.07 42.66 76.42 / 74.54 76.12 / 74.22 0.918 0.651 46.27 47.99 79.93 / 79.66 79.08 / 77.68 0.642 0.653
CENET 38.00 42.32 77.49 / 74.64 77.35 / 74.28 0.916 0.654 50.72 51.85 81.46 / 81.62 80.78 / 81.17 0.590 0.698
TETFN 35.81 41.79 77.49 / 75.60 77.35 / 75.35 0.910 0.657 35.81 52.35 77.49 / 75.60 77.35 / 75.35 0.910 0.657
TFR-Net 34.70 40.13 74.70 / 73.52 73.57 / 72.70 0.987 0.622 51.04 40.13 80.47 / 81.61 79.29 / 79.99 0.604 0.672
ALMT 35.33 40.33 77.64 / 75.70 77.94 / 76.24 0.927 0.645 46.61 47.82 82.71 / 81.65 82.82 / 81.83 0.607 0.669
LNLN 39.74 45.14 79.22 / 76.87 79.53 / 77.34 0.891 0.668 48.75 49.95 81.70 / 81.68 81.95 / 81.89 0.616 0.677
Random Missing Rate $r=0.3$
MISA 34.60 38.97 75.76 / 74.54 75.82 / 74.51 0.989 0.618 43.99 43.40 73.32 / 77.28 68.91 / 72.25 0.724 0.615
Self-MM 34.89 40.67 76.37 / 74.98 75.68 / 74.94 0.967 0.614 47.23 48.07 77.63 / 79.99 75.69 / 77.74 0.653 0.610
MMIM 36.83 40.43 74.08 / 71.91 73.47 / 71.28 0.974 0.612 43.25 44.73 77.08 / 77.79 75.46 / 74.49 0.690 0.597
CENET 34.74 38.97 76.83 / 72.01 76.56 / 71.30 0.983 0.605 48.49 49.37 78.65 / 80.02 77.34 / 78.94 0.636 0.640
TETFN 33.24 38.58 75.25 / 73.42 74.77 / 72.78 0.982 0.607 33.24 50.23 75.25 / 73.42 74.77 / 72.78 0.982 0.607
TFR-Net 32.55 38.34 72.36 / 71.28 70.12 / 69.58 1.065 0.572 48.75 38.34 77.48 / 79.29 75.43 / 76.52 0.650 0.604
ALMT 33.04 37.17 75.15 / 72.94 75.51 / 73.66 0.992 0.596 43.04 44.05 80.94 / 79.94 81.15 / 80.20 0.632 0.598
LNLN 38.00 42.81 77.29 / 75.46 77.56 / 75.68 0.953 0.617 47.36 48.40 80.11 / 80.45 80.44 / 80.91 0.648 0.629
Random Missing Rate $r=0.4$
MISA 32.65 35.37 73.88 / 72.59 73.88 / 72.49 1.041 0.585 40.87 39.53 70.46 / 75.04 64.02 / 67.93 0.780 0.561
Self-MM 31.20 36.30 73.17 / 71.96 71.74 / 71.75 1.027 0.579 44.40 45.04 75.02 / 78.09 72.01 / 74.48 0.694 0.554
MMIM 33.38 35.76 70.84 / 68.90 69.69 / 67.80 1.034 0.576 40.84 41.86 74.56 / 76.15 71.98 / 71.40 0.732 0.542
CENET 32.26 36.15 73.38 / 71.53 72.75 / 70.26 1.031 0.574 47.12 47.74 76.03 / 78.57 73.87 / 76.75 0.678 0.587
TETFN 30.66 35.28 72.05 / 70.07 70.79 / 68.58 1.051 0.571 30.66 47.62 72.05 / 70.07 70.79 / 68.58 1.051 0.571
TFR-Net 30.17 35.76 68.75 / 67.74 64.71 / 64.41 1.142 0.537 46.70 35.76 74.74 / 77.65 71.67 / 73.71 0.688 0.548
ALMT 31.44 35.03 73.12 / 71.14 73.85 / 72.47 1.045 0.56 40.40 41.21 79.40 / 79.16 79.68 / 79.50 0.651 0.536
LNLN 36.49 41.11 76.01 / 74.25 76.31 / 74.67 0.987 0.594 45.99 46.88 78.49 / 79.70 78.98 / 80.46 0.673 0.592
Random Missing Rate $r=0.5$
MISA 28.14 30.61 70.53 / 69.34 70.50 / 69.20 1.124 0.519 38.12 36.05 67.38 / 73.21 58.38 / 64.14 0.834 0.492
Self-MM 26.97 31.39 67.43 / 67.54 64.27 / 66.81 1.129 0.503 42.70 43.14 71.97 / 75.81 67.40 / 70.38 0.733 0.477
MMIM 28.23 29.89 68.09 / 66.52 66.15 / 64.59 1.128 0.501 38.68 39.21 71.75 / 74.45 67.70 / 67.96 0.775 0.470
CENET 28.33 30.90 72.46 / 66.08 71.10 / 63.50 1.130 0.496 45.12 45.52 73.33 / 77.16 69.80 / 74.14 0.720 0.515
TETFN 27.55 31.34 67.23 / 65.06 64.30 / 61.78 1.157 0.492 27.55 45.63 67.23 / 65.06 64.30 / 61.78 1.157 0.492
TFR-Net 25.85 30.71 64.83 / 63.02 58.04 / 56.64 1.270 0.443 45.00 30.71 71.53 / 75.69 66.88 / 70.07 0.730 0.471
ALMT 28.42 31.25 68.24 / 65.94 69.74 / 68.54 1.138 0.485 37.82 38.34 77.40 / 77.48 77.73 / 77.80 0.683 0.461
LNLN 33.92 38.39 73.37 / 71.86 73.70 / 72.30 1.059 0.536 44.90 45.59 76.44 / 78.10 77.23 / 79.30 0.710 0.529
Random Missing Rate $r=0.6$
MISA 24.68 27.12 66.97 / 65.84 66.94 / 65.69 1.200 0.441 36.16 33.30 65.55 / 72.30 54.64 / 62.12 0.875 0.415
Self-MM 24.34 27.31 63.47 / 63.36 58.94 / 62.07 1.209 0.425 41.47 41.75 69.33 / 73.93 63.01 / 66.76 0.762 0.401
MMIM 25.41 27.11 63.67 / 62.49 60.87 / 59.48 1.208 0.418 37.13 37.48 68.83 / 73.16 63.09 / 65.43 0.808 0.402
CENET 24.54 26.53 67.58 / 61.47 64.87 / 57.86 1.215 0.415 44.45 44.64 70.50 / 75.39 65.27 / 70.86 0.749 0.446
TETFN 25.12 27.99 63.42 / 61.23 58.68 / 56.08 1.238 0.417 25.12 44.07 63.42 / 61.23 58.68 / 56.08 1.238 0.417
TFR-Net 24.05 28.33 61.64 / 59.47 52.44 / 50.53 1.371 0.363 43.88 28.33 68.80 / 74.05 62.51 / 67.07 0.762 0.397
ALMT 25.41 27.36 64.53 / 62.15 66.81 / 65.87 1.214 0.407 35.99 36.30 74.98 / 76.26 75.44 / 76.71 0.710 0.395
LNLN 30.37 34.35 69.00 / 67.69 69.19 / 67.99 1.147 0.458 43.52 44.00 73.82 / 76.50 75.03 / 78.33 0.736 0.471
Random Missing Rate $r=0.7$
MISA 21.14 23.27 65.09 / 63.89 65.07 / 63.74 1.257 0.381 34.54 31.21 64.28 / 71.71 51.82 / 60.65 0.906 0.344
Self-MM 20.70 23.81 61.74 / 61.46 55.11 / 58.97 1.271 0.339 39.93 40.12 66.79 / 72.55 58.05 / 63.45 0.786 0.329
MMIM 22.35 24.00 61.23 / 59.18 57.15 / 54.36 1.267 0.342 35.25 35.47 66.89 / 72.26 58.90 / 63.26 0.834 0.341
CENET 22.35 59.86 63.82 / 59.43 53.79 / 54.22 1.269 0.335 43.93 44.03 67.50 / 73.39 59.88 / 67.02 0.776 0.384
TETFN 23.13 25.27 61.13 / 58.65 53.79 / 50.77 1.293 0.337 23.13 43.15 61.13 / 58.65 53.79 / 50.77 1.293 0.337
TFR-Net 23.71 26.92 59.91 / 57.34 48.41 / 45.48 1.454 0.276 42.91 26.92 66.64 / 72.77 58.32 / 64.02 0.786 0.322
ALMT 23.71 24.97 61.84 / 59.67 65.30 / 65.19 1.266 0.336 34.78 34.95 71.62 / 73.98 72.24 / 74.54 0.743 0.315
LNLN 27.79 31.19 65.95 / 65.01 65.95 / 65.14 1.219 0.383 42.22 42.56 71.55 / 74.74 73.49 / 77.40 0.762 0.408
Random Missing Rate $r=0.8$
MISA 19.92 20.99 63.56 / 62.24 63.16 / 61.67 1.311 0.321 33.29 29.51 63.43 / 71.30 49.95 / 59.69 0.927 0.267
Self-MM 19.29 22.11 59.55 / 58.26 49.98 / 53.56 1.313 0.282 38.69 38.78 65.07 / 71.83 54.44 / 61.49 0.805 0.259
MMIM 20.26 21.77 58.33 / 55.30 52.46 / 47.89 1.312 0.287 33.64 33.71 64.97 / 71.57 54.76 / 61.45 0.858 0.269
CENET 21.14 21.67 60.93 / 57.53 54.68 / 50.80 1.314 0.274 42.71 42.74 65.88 / 72.16 56.80 / 64.67 0.798 0.316
TETFN 22.01 23.76 59.40 / 56.85 48.73 / 45.59 1.337 0.274 22.01 42.11 59.40 / 56.85 48.73 / 45.59 1.337 0.274
TFR-Net 23.23 27.70 58.49 / 55.98 44.70 / 41.88 1.497 0.155 42.23 27.70 65.05 / 71.95 54.91 / 61.82 0.807 0.241
ALMT 23.13 23.66 60.37 / 58.31 65.45 / 66.14 1.310 0.273 34.01 34.09 68.15 / 71.48 69.12 / 72.28 0.774 0.231
LNLN 26.34 28.23 62.75 / 62.10 62.56 / 62.03 1.283 0.314 40.76 40.97 68.62 / 72.86 71.83 / 76.80 0.791 0.325
Random Missing Rate $r=0.9$
MISA 17.78 18.41 58.64 / 58.21 56.84 / 56.19 1.369 0.226 32.29 28.03 62.95 / 71.07 48.80 / 59.12 0.941 0.180
Self-MM 18.32 19.78 58.59 / 55.25 46.16 / 47.46 1.353 0.197 37.46 37.50 63.85 / 71.24 51.32 / 59.72 0.821 0.188
MMIM 18.95 19.53 55.29 / 51.65 47.33 / 40.89 1.357 0.186 32.61 32.67 63.69 / 71.10 51.26 / 59.99 0.877 0.197
CENET 19.15 19.10 58.99 / 54.76 50.01 / 46.58 1.357 0.181 42.08 42.08 64.14 / 70.42 54.27 / 62.33 0.814 0.254
TETFN 20.75 21.19 58.43 / 55.88 45.24 / 42.12 1.378 0.186 20.75 41.55 58.43 / 55.88 45.24 / 42.12 1.378 0.186
TFR-Net 21.67 25.12 57.93 / 55.44 43.01 / 40.18 1.534 0.155 41.73 25.12 63.64 / 71.34 52.02 / 59.99 0.820 0.175
ALMT 20.31 20.50 57.32 / 56.66 64.92 / 67.82 1.349 0.205 34.40 34.40 61.41 / 68.65 63.32 / 69.83 0.810 0.138
LNLN 22.98 23.86 56.50 / 56.51 56.32 / 56.47 1.349 0.202 40.10 40.19 64.83 / 71.51 70.60 / 77.52 0.820 0.221
Table 10: Details of Robust Comparison on SIMS with different random missing rates.
Method Acc-5 Acc-3 Acc-2 F1 MAE Corr
Random Missing Rate $r=0$
MISA 40.55 63.38 78.19 77.22 0.449 0.576
Self-MM 40.77 64.92 78.26 78.00 0.421 0.584
MMIM 37.42 60.69 75.42 73.10 0.475 0.528
CENET 23.85 54.05 68.71 57.82 0.578 0.137
TETFN 41.94 65.86 80.23 79.25 0.424 0.589
TFR-Net 33.85 54.12 69.15 58.44 0.562 0.254
ALMT 23.41 54.78 75.64 76.27 0.527 0.536
LNLN 38.66 63.97 75.93 79.89 0.458 0.570
Random Missing Rate $r=0.1$
MISA 38.88 63.02 77.39 75.82 0.461 0.561
Self-MM 40.26 63.53 77.32 76.76 0.433 0.563
MMIM 37.27 60.90 74.25 72.08 0.473 0.529
CENET 22.83 53.98 68.57 57.36 0.580 0.136
TETFN 41.36 64.62 78.92 77.70 0.432 0.578
TFR-Net 30.12 53.25 68.85 59.38 0.596 0.203
ALMT 22.10 55.14 74.40 75.19 0.530 0.537
LNLN 38.51 62.73 76.29 80.07 0.458 0.562
Random Missing Rate $r=0.2$
MISA 38.15 59.23 74.33 71.70 0.489 0.490
Self-MM 38.37 61.71 74.98 73.71 0.464 0.500
MMIM 37.27 57.33 72.36 69.80 0.504 0.460
CENET 22.25 54.20 68.57 57.64 0.583 0.132
TETFN 39.46 61.56 75.49 73.59 0.457 0.527
TFR-Net 29.03 53.61 68.64 59.74 0.619 0.191
ALMT 21.08 53.17 72.65 73.90 0.541 0.485
LNLN 38.88 61.78 74.76 78.53 0.474 0.513
Random Missing Rate $r=0.3$
MISA 36.40 59.30 74.11 70.40 0.505 0.464
Self-MM 37.93 59.81 74.76 72.85 0.474 0.487
MMIM 37.71 58.06 72.36 69.52 0.512 0.436
CENET 21.44 54.05 68.42 57.41 0.578 0.175
TETFN 38.80 61.92 75.86 73.28 0.463 0.521
TFR-Net 27.64 52.30 68.42 59.88 0.640 0.182
ALMT 20.35 50.62 72.06 73.64 0.546 0.469
LNLN 38.37 60.98 74.25 78.60 0.478 0.509
Random Missing Rate $r=0.4$
MISA 34.86 57.33 72.87 67.52 0.523 0.436
Self-MM 34.57 58.28 73.30 70.36 0.482 0.479
MMIM 34.57 55.36 69.95 66.49 0.533 0.399
CENET 22.54 54.12 68.49 57.68 0.583 0.141
TETFN 35.81 58.93 73.81 70.66 0.473 0.504
TFR-Net 25.31 51.86 67.91 59.16 0.664 0.176
ALMT 19.91 49.45 70.75 72.97 0.549 0.470
LNLN 37.49 60.03 73.81 78.82 0.491 0.481
Method Acc-5 Acc-3 Acc-2 F1 MAE Corr
Random Missing Rate $r=0.5$
MISA 30.56 54.78 71.26 64.16 0.552 0.367
Self-MM 32.02 53.90 71.41 67.11 0.517 0.390
MMIM 33.41 52.37 68.49 64.81 0.553 0.336
CENET 23.12 54.05 68.71 57.92 0.588 0.107
TETFN 33.48 56.24 72.43 67.30 0.512 0.394
TFR-Net 24.65 52.37 67.47 58.66 0.685 0.171
ALMT 18.38 47.12 68.27 71.22 0.563 0.395
LNLN 36.40 57.70 72.72 78.77 0.513 0.412
Random Missing Rate $r=0.6$
MISA 27.72 53.97 70.46 61.81 0.578 0.286
Self-MM 29.10 51.86 70.02 64.21 0.548 0.313
MMIM 29.18 49.31 67.91 63.86 0.578 0.270
CENET 22.46 53.69 69.00 58.64 0.592 0.102
TETFN 29.90 53.54 70.97 64.19 0.545 0.309
TFR-Net 24.80 52.59 67.03 58.30 0.696 0.157
ALMT 18.67 43.69 66.81 70.69 0.574 0.322
LNLN 33.70 54.63 71.55 79.56 0.535 0.352
Random Missing Rate $r=0.7$
MISA 24.87 52.52 69.95 59.54 0.601 0.167
Self-MM 25.53 50.62 69.58 62.28 0.571 0.198
MMIM 28.59 46.53 66.89 62.23 0.595 0.190
CENET 21.81 53.32 67.69 57.87 0.599 0.070
TETFN 27.86 51.06 69.29 61.09 0.572 0.190
TFR-Net 23.78 52.30 67.18 58.15 0.707 0.163
ALMT 18.02 38.66 65.57 70.27 0.586 0.218
LNLN 30.27 51.64 69.73 79.37 0.558 0.261
Random Missing Rate $r=0.8$
MISA 22.69 52.22 69.37 57.82 0.610 0.092
Self-MM 22.03 50.77 69.51 60.68 0.585 0.138
MMIM 22.32 44.35 65.28 60.53 0.607 0.145
CENET 21.73 52.15 67.47 58.44 0.599 0.074
TETFN 23.48 49.60 69.88 60.92 0.584 0.154
TFR-Net 22.97 52.74 67.54 57.55 0.721 0.100
ALMT 18.60 34.06 64.19 69.64 0.597 0.133
LNLN 27.94 50.47 69.58 80.23 0.580 0.183
Random Missing Rate $r=0.9$
MISA 20.64 52.95 69.22 57.01 0.617 0.041
Self-MM 22.17 52.15 68.92 58.32 0.586 0.111
MMIM 20.35 42.67 65.72 59.64 0.610 0.096
CENET 20.86 48.07 65.72 58.18 0.609 -0.002
TETFN 22.10 45.73 68.92 58.75 0.590 0.108
TFR-Net 23.05 53.76 69.08 57.71 0.721 0.088
ALMT 19.47 26.91 66.23 73.76 0.596 0.076
LNLN 26.19 47.48 68.64 80.42 0.591 0.127
B.5 Visualization of Confusion Matrix
To verify the effectiveness of the method, we visualized the confusion matrix of several representative methods on the MOSI, MOSEI, and SIMS datasets in Figure 5, Figure 6, and Figure 7, respectively. Obviously, as the missing rate $r$ increases, the predictions of all methods on all datasets tend to favor certain categories. This phenomenon indicates that the models hardly learn useful knowledge for prediction but instead ensure high accuracy by being lazy. For example, on the MOSI dataset, Self-MM predicts all samples as weak negative when the missing rate is 0.9. Although the accuracy of Self-MM is high when the missing rate $r$ is high, it does not demonstrate that the model is more robust.
In addition, TETFN and TFR-Net tend to randomly predict within the negative and weak negative categories under high noise, indicating that their lazy behavior is not as severe as methods like Self-MM, MMIM, and MISA. Unlike other methods, although LNLN also suffers from the impact of data missing, the model’s lazy behavior is less severe. For example, at a missing rate of 0.9, the model is still attempting to predict the samples, and there is no case of predictions favoring a particular category to an extreme extent.
From Figure 6 and Figure 7, we can see that a similar phenomenon also occurs in the MOSEI and SIMS datasets. These observations demonstrate that methods targeting complete data often fail when modeling highly incomplete data. Therefore, the design of methods specifically for modeling incomplete data is necessary.
Figure 5: Seven-category Confusion matrix of several representative methods on MOSI dataset. Note: 0-6 denote strongly negative, weakly negative, negative, neutral, weakly positive, positive, and strongly positive, respectively.
Figure 6: Seven-category Confusion matrix of several representative methods on MOSEI dataset. Note: 0-6 denote strongly negative, weakly negative, negative, neutral, weakly positive, positive, and strongly positive, respectively.
Figure 7: Five-category Confusion matrix of several representative methods on SIMS dataset. Note: 0-4 denote weakly negative, negative, neutral, weakly positive, positive, respectively.
B.6 Visualization of Representations
Figure 8 shows the language feature $H^{1}_{l}$ extracted without applying data missing and the corrected dominant feature $H^{1}_{d}$ with different data missing rates. we can see that the model tends to learn $H^{1}_{d}$ that is consistent with the distribution of the $H^{1}_{l}$, indicating that LNLN can complement the information of the language modality when data is missing. Additionally, when the missing rate $r$ is 1.0 (no valid information for the input), it becomes difficult for the model to learn $H^{1}_{d}$ that is consistent with the language feature. This also suggests that the generator is indeed trying to learn $H^{1}_{d}$ for fusion in DMML, rather than just guessing.
Figure 8: Visualization of the corrected dominant feature $H^{1}_{d}$ with different data missing rate and the language feature $H^{1}_{l}$ without data missing.
B.7 Analysis of the generalization to modality missing scenarios
To explore the generalization of the method, we further evaluated it in the special case of data missing, i.e., modality missing scenarios. The overall performance of the model is shown in Table 11 and Table 12. The detailed performance of the model is shown in Table 13 and Table 14. It is evident that LNLN, ALMT, TETFN, and Self-MM all exhibit excellent generalization.
We believe that the Dominant Modality Correction (DMC) proposed in LNLN, the Adaptive Hyper-modality Learning (AHL) proposed in ALMT, and the Unimodal Label Generation Module (ULGM) proposed/used in Self-MM and TETFN deserve more research attention in the future. These modules and their corresponding ideas have significant reference value for achieving robust Multimodal Sentiment Analysis (MSA).
However, we also found that many methods converge to the same value when the noise is high. This is similar to the phenomenon observed in previous data missing scenarios, where many models exhibited lazy behavior. Therefore, handling modality missing and improving model robustness in different noise scenarios remain critical challenges that need to be addressed in the field of MSA.
Table 11: Generalization comparison of the overall performance on MOSI and MOSEI with random modality missing. Note: the parameters used for evaluation are consistent with those used for testing in random data missing. The smaller MAE indicate the better performance.
Method MOSI MOSEI
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-7 Acc-5 Acc-2 F1 MAE Corr
MISA 29.44 31.76 67.64 / 67.27 64.81 / 64.30 1.096 0.461 41.45 40.32 73.94 / 77.49 66.65 / 71.30 0.752 0.438
Self-MM 29.74 36.40 71.45 / 68.17 63.71 / 61.91 1.052 0.450 45.17 46.05 74.12 / 77.85 67.00 / 71.81 0.684 0.447
MMIM 31.28 33.74 68.34 / 65.12 63.38 / 58.28 1.058 0.444 41.36 42.53 73.19 / 76.14 68.11 / 70.18 0.732 0.448
CENET 30.30 33.94 70.37 / 66.67 62.66 / 58.61 1.068 0.450 47.18 47.83 74.67 / 77.34 70.68 / 74.08 0.685 0.535
TETFN 32.66 36.26 70.17 / 68.15 62.47 / 60.18 1.061 0.446 47.41 48.41 73.87 / 77.43 66.70 / 71.45 0.684 0.480
TFR-Net 29.76 34.96 70.49 / 67.95 62.87 / 60.08 1.177 0.451 47.33 48.08 73.50 / 77.45 66.89 / 71.28 0.700 0.427
ALMT 27.87 30.73 70.53 / 67.47 75.89 / 73.35 1.130 0.447 25.94 27.46 60.66 / 63.37 70.63 / 71.24 0.711 0.354
LNLN 31.93 34.86 68.53 / 65.64 72.01 / 69.65 1.093 0.423 43.85 44.50 73.52 / 77.04 80.85 / 83.19 0.733 0.425
Table 12: Generalization comparison of the overall performance on SIMS dataset with random modality missing. Note: the parameters used for evaluation are consistent with those used for testing in random data missing. The smaller MAE indicate the better performance.
Method Acc-5 Acc-3 Acc-2 F1 MAE Corr
MISA 29.62 56.82 73.61 66.32 0.535 0.330
Self-MM 32.49 56.36 74.30 80.59 0.522 0.289
MMIM 27.21 50.79 71.52 65.48 0.546 0.286
CENET 19.95 43.63 64.08 56.96 0.601 0.033
TETFN 30.61 50.98 74.77 68.00 0.514 0.317
TFR-Net 24.69 52.57 69.31 58.03 0.629 0.133
ALMT 21.82 39.39 72.30 78.86 0.563 0.287
LNLN 30.39 54.32 72.67 80.87 0.527 0.287
Table 13: Generalization comparison on MOSI and MOSEI with random modality missing. Note: The smaller MAE indicate the better performance.
Method MOSI MOSEI
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-7 Acc-5 Acc-2 F1 MAE Corr
Language Modality Missing
MISA 15.60 15.65 52.18 / 53.74 48.09 / 49.42 1.422 0.171 30.93 26.75 71.02 / 62.85 58.99 / 48.51 0.946 0.128
Self-MM 16.86 19.29 57.77 / 54.37 42.31 / 43.81 1.381 0.129 36.59 36.59 71.02 / 63.03 58.99 / 49.22 0.832 0.152
MMIM 17.44 17.20 54.67 / 49.61 46.95 / 37.63 1.363 0.153 31.96 31.97 62.71 / 71.01 54.70 / 59.43 0.886 0.195
CENET 17.54 17.54 57.67 / 51.80 42.26 / 35.71 1.387 0.118 41.70 41.70 63.23 / 69.38 52.41 / 60.67 0.821 0.238
TETFN 21.19 21.14 57.77 / 55.25 42.31 / 39.32 1.403 0.097 41.31 41.31 62.82 / 71.01 48.53 / 58.98 0.831 0.148
TFR-Net 23.13 25.95 57.82 / 55.30 42.43 / 39.43 1.563 0.167 41.40 41.43 62.85 / 71.02 48.88 / 58.99 0.829 0.163
ALMT 19.29 19.68 56.40 / 55.05 66.99 / 66.01 1.394 0.139 21.58 21.58 54.28 / 57.38 69.52 / 70.01 0.867 0.071
LNLN 18.80 17.68 52.18 / 49.03 58.89 / 56.84 1.427 0.075 39.10 39.10 62.85 / 71.02 77.19 / 83.06 0.847 0.156
Audio Modality Missing
MISA 43.49 48.44 83.74 / 81.58 83.72 / 81.49 0.762 0.776 51.59 53.91 84.94 / 83.91 84.62 / 83.46 0.561 0.759
Self-MM 42.81 52.33 85.22 / 83.24 85.19 / 83.26 0.720 0.790 53.91 55.66 85.31 / 84.70 85.06 / 84.64 0.532 0.763
MMIM 45.39 49.52 83.08 / 81.15 83.00 / 81.02 0.744 0.777 51.01 53.26 83.72 / 81.51 83.60 / 81.29 0.572 0.726
CENET 43.05 50.39 83.08 / 81.54 83.06 / 81.54 0.750 0.785 54.34 55.98 85.37 / 82.08 85.35 / 82.44 0.531 0.770
TETFN 44.07 51.31 82.52 / 81.00 82.57 / 80.99 0.719 0.794 54.15 55.94 85.08 / 83.99 85.05 / 84.05 0.543 0.753
TFR-Net 37.71 45.48 82.72 / 80.52 82.71 / 80.45 0.821 0.759 53.75 55.34 84.95 / 84.65 84.69 / 84.32 0.550 0.745
ALMT 36.49 41.84 84.60 / 79.88 84.81 / 80.70 0.865 0.767 30.51 33.55 67.13 / 70.05 71.76 / 72.59 0.535 0.556
LNLN 45.05 52.04 84.86 / 82.21 85.12 / 82.43 0.760 0.772 50.63 51.92 84.39 / 83.09 84.71 / 83.34 0.610 0.736
Visual Modality Missing
MISA 43.29 47.09 82.37 / 81.00 82.43 / 80.99 0.777 0.777 51.76 53.80 85.25 / 83.99 85.11 / 83.76 0.550 0.757
Self-MM 42.76 55.47 85.11 / 82.80 85.09 / 82.82 0.722 0.789 53.68 55.47 85.20 / 84.62 84.95 / 84.59 0.533 0.761
MMIM 44.75 49.85 83.54 / 81.83 83.52 / 81.78 0.740 0.778 50.45 52.86 83.71 / 81.31 83.53 / 81.09 0.579 0.722
CENET 43.20 50.39 83.08 / 81.49 83.06 / 81.48 0.748 0.785 52.95 54.48 84.99 / 81.04 85.03 / 81.56 0.544 0.761
TETFN 44.22 51.46 82.62 / 81.10 82.67 / 81.09 0.719 0.794 52.95 55.34 84.84 / 83.78 84.81 / 83.86 0.544 0.761
TFR-Net 37.85 45.87 83.54 / 81.10 83.41 / 80.95 0.799 0.760 52.94 54.45 83.97 / 83.48 83.93 / 83.18 0.577 0.726
ALMT 36.39 41.93 84.81 / 79.98 84.81 / 80.76 0.863 0.768 29.80 32.95 66.81 / 69.41 71.72 / 72.23 0.559 0.747
LNLN 45.05 51.99 84.91 / 82.26 85.17 / 82.48 0.759 0.772 50.16 51.44 84.09 / 83.41 84.43 / 83.76 0.577 0.729
Language & Audio Modality Missing
MISA 15.55 15.60 55.29 / 55.20 49.53 / 49.24 1.419 0.098 31.17 26.76 62.85 / 71.02 48.51 / 58.99 0.961 0.121
Self-MM 16.86 19.29 57.77 / 54.03 42.31 / 43.31 1.381 0.130 36.61 36.61 63.01 / 71.02 49.18 / 58.99 0.831 0.144
MMIM 17.59 17.40 52.24 / 48.20 39.57 / 31.53 1.389 0.024 31.91 31.92 62.31 / 70.94 54.32 / 59.28 0.886 0.182
CENET 17.40 17.40 57.67 / 51.85 42.26 / 35.65 1.387 0.110 41.40 41.40 60.69 / 64.21 51.22 / 56.68 0.823 0.236
TETFN 21.19 21.14 57.77 / 55.25 42.31 / 39.32 1.403 0.098 41.31 41.31 62.82 / 71.00 48.51 / 58.98 0.832 0.144
TFR-Net 21.38 23.96 58.03 / 55.59 43.37 / 40.54 1.554 0.107 41.37 41.38 62.81 / 71.02 48.75 / 58.99 0.829 0.161
ALMT 19.24 19.24 56.35 / 54.96 66.89 / 65.86 1.397 0.104 21.57 21.57 54.28 / 57.28 69.52 / 70.09 0.870 0.091
LNLN 18.80 17.68 52.18 / 49.03 58.89 / 56.84 1.427 0.072 34.72 34.72 62.85 / 71.02 77.19 / 83.06 0.900 0.145
Language & Visual Modality Missing
MISA 15.45 15.40 48.98 / 50.87 41.79 / 43.51 1.427 0.169 31.57 26.89 62.85 / 71.02 48.51 / 58.99 0.936 0.105
Self-MM 16.47 19.24 57.77 / 51.75 42.31 / 35.44 1.386 0.072 36.55 36.55 62.85 / 71.02 48.51 / 58.99 0.838 0.101
MMIM 17.93 17.40 53.05 / 48.64 43.80 / 36.52 1.366 0.155 32.12 32.12 62.88 / 70.94 48.91 / 59.00 0.891 0.143
CENET 17.54 17.54 57.67 / 51.80 42.26 / 35.71 1.387 0.118 41.36 41.36 61.78 / 67.49 51.87 / 59.63 0.838 0.112
TETFN 21.19 21.14 57.77 / 55.25 42.31 / 39.32 1.403 0.097 41.36 41.36 62.81 / 71.02 48.51 / 58.99 0.840 0.017
TFR-Net 22.50 25.07 57.77 / 55.25 42.31 / 39.32 1.486 0.154 41.36 41.36 62.47 / 71.02 51.16 / 58.99 0.839 0.039
ALMT 19.29 19.77 56.45 / 55.10 67.09 / 66.11 1.394 0.136 21.54 21.54 54.28 / 57.01 69.52 / 70.35 0.874 -0.086
LNLN 18.80 17.68 52.18 / 49.03 58.89 / 56.84 1.427 0.075 38.41 38.41 62.85 / 71.02 77.19 / 83.06 0.853 0.052
Audio & Visual Modality Missing
MISA 43.25 48.40 83.28 / 81.24 83.29 / 81.17 0.768 0.776 51.67 53.80 84.88 / 83.97 84.64 / 83.64 0.558 0.756
Self-MM 42.71 52.77 85.06 / 82.80 85.04 / 82.82 0.722 0.789 53.67 55.43 85.32 / 84.74 85.06 / 84.67 0.535 0.761
MMIM 44.56 51.07 83.48 / 81.29 83.42 / 81.17 0.748 0.777 50.70 53.07 83.79 / 81.14 83.61 / 80.96 0.579 0.722
CENET 43.05 50.39 83.08 / 81.54 83.06 / 81.54 0.750 0.785 52.71 54.22 84.64 / 80.66 84.71 / 81.23 0.545 0.760
TETFN 44.12 51.36 82.57 / 81.05 82.62 / 81.04 0.719 0.794 53.46 55.22 84.83 / 83.79 84.80 / 83.85 0.549 0.747
TFR-Net 36.00 43.44 83.08 / 79.98 83.01 / 79.82 0.838 0.758 53.13 54.54 83.97 / 83.49 83.91 / 83.18 0.578 0.726
ALMT 36.49 41.93 84.55 / 79.83 84.76 / 80.65 0.864 0.767 30.64 33.57 67.14 / 69.11 71.77 / 72.15 0.560 0.748
LNLN 45.10 52.04 84.86 / 82.26 85.12 / 82.48 0.760 0.772 50.10 51.38 84.10 / 82.69 84.39 / 82.86 0.609 0.730
Table 14: Generalization comparison on SIMS with random modality missing. Note: The smaller MAE indicate the better performance.
Method Acc-5 Acc-3 Acc-2 F1 MAE Corr
Language Modality Missing
MISA 18.60 50.77 69.23 56.75 0.595 0.104
Self-MM 17.22 41.28 67.76 57.76 0.617 0.055
MMIM 17.22 41.28 67.76 57.76 0.617 0.055
CENET 20.71 49.24 68.20 57.67 0.601 0.037
TETFN 19.26 36.11 69.37 56.82 0.603 0.045
TFR-Net 22.25 54.27 69.30 56.78 0.712 0.008
ALMT 21.15 24.95 69.37 81.91 0.595 0.049
LNLN 21.37 45.30 69.37 81.91 0.597 0.005
Audio Modality Missing
MISA 38.37 61.92 77.68 75.17 0.466 0.574
Self-MM 40.92 64.99 77.32 76.76 0.421 0.585
MMIM 37.27 60.61 75.20 72.96 0.475 0.527
CENET 21.59 53.90 68.93 57.45 0.582 0.141
TETFN 41.94 65.86 80.16 79.16 0.424 0.589
TFR-Net 26.40 52.45 69.66 59.21 0.569 0.239
ALMT 23.05 59.45 75.27 76.36 0.531 0.529
LNLN 39.53 64.19 75.78 79.69 0.454 0.570
Viusal Modality Missing
MISA 40.84 63.67 78.19 77.07 0.458 0.574
Self-MM 41.21 64.99 74.98 73.71 0.421 0.584
MMIM 37.34 60.47 75.57 73.43 0.475 0.528
CENET 19.70 38.51 59.08 56.38 0.600 0.027
TETFN 41.94 65.86 80.16 79.19 0.424 0.589
TFR-Net 30.20 54.56 69.15 58.86 0.566 0.247
ALMT 22.83 55.51 75.35 75.75 0.529 0.535
LNLN 38.80 63.31 76.22 80.11 0.458 0.569
Method Acc-5 Acc-3 Acc-2 F1 MAE Corr
Language & Audio Modalities Missing
MISA 21.23 54.27 69.37 56.82 0.613 0.064
Self-MM 19.69 54.20 74.76 72.85 0.594 0.059
MMIM 17.22 41.28 67.91 57.86 0.617 0.026
CENET 20.42 49.60 66.31 57.49 0.601 0.002
TETFN 19.26 36.11 69.37 56.82 0.603 0.04
TFR-Net 21.96 46.98 69.44 57.12 0.635 0.041
ALMT 21.30 16.05 69.37 81.91 0.594 0.039
LNLN 21.37 43.25 69.37 81.91 0.598 -0.008
Language & Visual Modalities Missing
MISA 19.70 48.58 69.37 56.95 0.600 0.091
Self-MM 19.48 53.98 73.30 70.36 0.594 0.056
MMIM 17.22 40.99 67.40 57.70 0.617 0.055
CENET 18.45 33.55 65.50 59.55 0.612 -0.054
TETFN 19.33 36.11 69.37 56.82 0.603 0.045
TFR-Net 19.55 54.27 69.37 56.82 0.720 0.027
ALMT 19.99 20.57 69.15 81.08 0.596 0.043
LNLN 21.44 46.39 69.37 81.91 0.600 0.014
Audio & Viusal Modalities Missing
MISA 38.95 61.70 77.83 75.15 0.478 0.573
Self-MM 41.06 64.99 71.41 67.11 0.421 0.585
MMIM 36.98 60.10 75.27 73.18 0.476 0.527
CENET 18.82 36.98 56.46 53.23 0.608 0.045
TETFN 41.94 65.86 80.16 79.19 0.424 0.589
TFR-Net 27.79 52.88 68.93 59.40 0.571 0.234
ALMT 22.61 59.81 75.27 76.17 0.532 0.529
LNLN 39.82 63.46 75.93 79.72 0.454 0.569
Appendix C Social Impacts
Our proposed LNLN has a wide range of applications in real-world scenarios, such as healthcare and human-computer interaction. However, it might also be misused to monitor individuals based on their affections for illegal purposes, potentially posing negative social impacts.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope. They clearly outline the capabilities of the proposed LNLN and its performance in handling data missing scenarios, as well as emphasizing the comprehensive evaluation of existing methods and their significance in improving robust Multimodal Sentiment Analysis.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discussed the limitations of the work, which is specifically in the appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The article does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper reports the network structure required for reproduction as well as the hyperparameter settings. The code will be made public on Github after being accepted.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The datasets used in the thesis are open datasets and can be accessed by anyone upon request. In addition the code will be open-sourced at Github later.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Optimizer selection, data partitioning, and other details are described. Hyperparameters are chosen empirically.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: As shown in 8, we report the variance/fluctuation of several representative methods on the mosi dataset. It shows the overall stability of the different methods in the face of noise impacts of different levels.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report this information in Section Implementation Details.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This study meets the requirements of moral and ethical norms.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We describe this in the appendix.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The used data has been cited in the paper. The code will be available on Github after sort out.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [No]
  64. Justification: The dataset used is a public dataset, and the code will be available on Github after sort out.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve above problems.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This study does not involve the above issues and the datasets used are all publicly available datasets.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.